OpenStudy (vishweshshrimali5):

Find the real point (x,y) satisfying \(\large{3x^2 + 3y^2 - 4xy + 10x - 10y + 10 = 0}\)

OpenStudy (vishweshshrimali5):

@mukushla @ganeshie8 @Miracrown @BSwan

OpenStudy (vishweshshrimali5):

@dan815 @iambatman

mathslover (mathslover):

@mathslover

OpenStudy (vishweshshrimali5):

Sorry I forgot the most important one: @mathslover . Pls give a huge round of applause for him. :P

Miracrown (miracrown):

This looks like a rotated ellipse, so there's probably many points. Thinking... <.< >.>

OpenStudy (vishweshshrimali5):

One more point: do not use wolframalpha. Coz I am using it currently for this and I do not want to increase traffic on the side.

OpenStudy (vishweshshrimali5):

Hey @mathslover remember this discussion is being observed by a mod @mukushla . Lets be serious ;)

mathslover (mathslover):

lol Yeah :-) After all, I'm a lot famous! And sorry for those jokes. I will be serious now. (deleted my comments)

OpenStudy (vishweshshrimali5):

It is an ellipse @Miracrown

OpenStudy (anonymous):

haha :D

OpenStudy (vishweshshrimali5):

Important point: Do not look for its geometry or you are going to stuck very prettily (I speak from experience *sigh*). Look for algebra and ask for its help (a better word would be request algebra for its help).

OpenStudy (anonymous):

let's factor it

OpenStudy (vishweshshrimali5):

Okay let me present my side of solution:

Miracrown (miracrown):

3y^2 - (10+4x)y + (3x^2+10x+10) = 0 we can find a real point y = (10+4x) +- Sqrt[(10+4x)^2 - 4(3)(3x^2+10x+10)] all over 6 y = (10+4x) +- Sqrt[(10+4x)^2 - 4(3)(3x^2+10x+10)] all over 6 -4(3)(3x^2+10x+10) = -36x^2 -120x -120 so y = 1/6 * (10 + 4x +- Sqrt[-20-20x^2-40x]) Factor the 20 y = 1/3 (5+2x +- Sqrt[5] Sqrt[-1-2x-x^2]) so then we want -1-2x-x^2 non-negative best we can do is zero solving -x^2-2x-1 = 0 So the problem asks us for a real point If the thing under the square root is negative, we get a complex number

Miracrown (miracrown):

Therefore, we need it to be non-negative. But it's clearly negative for almost all x, so it's going to turn out that one when -x^2-2x-1 is zero do we get a real solution.

mathslover (mathslover):

I will prefer you to first make the coefficients of \(x^2\) and \(y^2\) as 1. i.e. divide by 3 both sides. And then factor.

Miracrown (miracrown):

this happens when x = -1 right? -x^2-2x-1 = - (x+1)^2 and when x = -1

OpenStudy (vishweshshrimali5):

Great work @Miracrown You figured it again. The corresponding value of y is 1.

Miracrown (miracrown):

and so y = 1/3 (5 + 2(-1)) = 1 (-1,1)

Miracrown (miracrown):

thank you

Miracrown (miracrown):

Right, so, not the easiest problem in the world, but also not the most difficult

OpenStudy (vishweshshrimali5):

Let me present an alternative method: I am going to simplify the LHS: \[3(x^2 + 2x + 1) + 3(y^2 - 2y +1) + 4(1+x-y-xy)\] \[= 3(x+1)^2 + 3(1-y)^2 + 4(1+x)(1-y)\] \[=(1+x)^2 + (1-y)^2 + 2[(1+x)^2 + (1-y)^2 + 2(1+x)(1-y)]\] \[=(1+x)^2 + (1-y)^2 + 2[(1+x)+(1-y)]^2\] Now this must be equal to 0. So, each term must be zero. So, x = -1 and y = 1

OpenStudy (vishweshshrimali5):

There you have it : another easy method.

OpenStudy (vishweshshrimali5):

Hint for you all: whenever you see such an equation look for representing it in the form of sum of squares and then equate each terms as 0.

OpenStudy (anonymous):

ROGER :)

OpenStudy (vishweshshrimali5):

Bravo !!

mathslover (mathslover):

Is there any way to go on from here? or is it completely the wrong method? \[ \left[ x + \cfrac{5}{3} \right] ^2 + \left [ y - \cfrac{5}{3} \right]^2 = \cfrac{20}{9} - \cfrac{4}{3} xy \\ \left[ x + \cfrac{5}{3} \right]^2 + \left[ y - \cfrac{5}{3} \right] ^2 + \cfrac{4}{3} xy = \cfrac{20}{9} \]

OpenStudy (vishweshshrimali5):

@mathslover Actually its not wrong as per. But, it is not going to help you much.

Miracrown (miracrown):

@vishweshshrimali5 Right, so what you're basically doing here is expressing the equation as a polynomial in two variables. And because equal power of each variable is linearly independent they all must be zero in order to have a root. I do not like this method because it's not quite as reliable, I think. Getting into this form (if always possible) is not so obvious. :P

mathslover (mathslover):

Oh.. fine! And yes, I do think that I did a mistake there. I will check that out, and see if it leads to something.

Miracrown (miracrown):

equal=each power* Haha, sorry, opinions!

OpenStudy (vishweshshrimali5):

Hey @Miracrown , opinions are always welcomed by me and I really appreciate them. Actually what I meant there was that its just another trick which may work under special conditions (which in math olympiads in fairly general). Its not a basic method but an intelligent one. Actually this question was given to us in a question bank. So, I thought to share it. I am going to share some more. You are all welcome to try them as well.

Miracrown (miracrown):

Haha, no problem. Yes, indeed Are you studying for a math olympiad or something?

OpenStudy (vishweshshrimali5):

Well I have already been trained for math olympiads in past. I have been regional topper for 2 consecutive years in KVS (our school organisation). Moreover, I have represented KVS in INMO (national level math olympiad). But, due to less time and more work, I couldn't perform well in it. Now I just go through the problems and try solving them even though they are not going to benefit me in my future engineering degree.

Miracrown (miracrown):

O.O you never know... much of the complex engineering problems at the undergraduate level will feel like solving some sort of 19th century math stumper ;)

OpenStudy (vishweshshrimali5):

Well I will love to go through those 19th century stumpers. :D

Miracrown (miracrown):

:)

mathslover (mathslover):

I just made that problem greatly complexed. \( 7(x^2 + y^2)^2 + 16 x^2 y^2 - 56 (x^2 + y^2) (xy) = 100 ( 1 + x^2 + y^2 -2xy + 2\sqrt{ (t - 2xy) } ) \) No one will ever like to solve this. I will better go with your method instead of my nasty one. :P Great work @vishweshshrimali5 and @Miracrown

OpenStudy (vishweshshrimali5):

Well its good to see that you at least tried.

mathslover (mathslover):

Yeah.. you never know, how much substitution I did in between. :( Though, this led me to learn some new things.

Miracrown (miracrown):

It's more important to fail than to succeed.